OpenStudy (anonymous):

Find the coordinates of the point Q on the curve -y=x^2-4x+7 at which the tangent is parallel to the line y+3x=9.

OpenStudy (anonymous):

yes EXP, anything to say ?

OpenStudy (anonymous):

-y=x^2-4x+7 or this is y=-x^2+4x-7 and this line is parallel to y+3x=9.or y=-3x+9.your slope here is m=-3 also the slope of y=-x^2+4x-7 is dy/dx=m=-2x-4 since the line is parallel therefore their slope are equal so m=m=-2x+4=-3 -2x=-7 x=7/2 sub this to the line eq to find y y=-x^2+4x-7 y=-(7/2)^2+4(7/2)-7 thus y=-21/4 now you coordinate Q(x,y)=Q(7/2,-21/4) ..ans...

OpenStudy (anonymous):

oopps typo that was dy/dx=m=-2x+4

OpenStudy (anonymous):

any question let me know

OpenStudy (anonymous):

Thank you so much. (: